# Quantum Energy Teleportation: A Novel Protocol Relaxing Previous Constraints

### v1

@@ Abstract
Quantum state teleportation (QST) is a foundational protocol for transferring quantum information across long distances. Building on this concept, quantum energy teleportation (QET) was introduced to facilitate the transfer of energy. However, QET has traditionally been thought to impose much stricter constraints than QST, including the need for the initial state to be the ground state of a specific interacting Hamiltonian, for Alice’s measurement to commute with the interaction terms, and for the ground state to exhibit entanglement. These constraints limit broader application of the QET protocol. In this work, we challenge these notions and demonstrate for the first time that none of these constraints are fundamentally necessary. We propose a novel QET protocol that relaxes all these constraints, offering a more flexible and generalized framework. Additionally, we introduce the concept of a “local effective Hamiltonian,” which removes the need for optimization techniques in determining Bob’s optimal energy extraction. These advancements enhance our understanding of QET and extends its broader applications for quantum technologies. To support our findings, we implement the protocol on quantum hardware, confirming its theoretical validity and experimental feasibility.
@@ I Introduction
In contrast, a related protocol, quantum energy teleportation (QET), was believed to enable the teleportation of energy Hotta (2009a, ). Similar to QST, QET relies on classical communication to ensure that energy transfer does not exceed the speed of light. However, QET imposes stricter constraints compared to QST. While QST only requires the sender and receiver to share an entangled state, QET further requires that this state be the ground state of an interacting Hamiltonian. Additionally, in QST, the sender performs a measurement and transmits the outcome to the receiver via a classical communication channel. In QET, however, the observable used for the measurement must also commute with the interaction term in the Hamiltonian, ensuring that the local energy of the receiver remains undisturbed prior to the teleportation process.
@@
Since its initial discovery in spin-chain systems Hotta (2008a, 2009a), QET has been extensively explored in many other physical systems. Theoretical studies have investigated its application in relativistic quantum field theory Hotta (2008b); Ikeda (2023a), trapped ions Hotta (2009b), harmonic oscillators Hotta (2010a), black-hole physics Hotta (2010b), linear harmonic chains Nambu and Hotta (2010), quantum Hall systems Yusa et al. (2011), Gibbs spin particles Frey et al. (2013); Trevison and Hotta (2015), squeezed vacuum states Hotta et al. (2014), topological orders Ikeda (2023b), and quantum information science Ikeda and Lowe (2024). More recently, experimental efforts have demonstrated progress toward realizing QET, including in an NMR system Rodríguez-Briones et al. (2023) and on superconducting quantum hardware Ikeda (2023c). The combination of QST and QET was proposed to enable a long-range QET protocol Ikeda (2023d). A trade-off relationship between QST and QET was proposed, suggesting their potential exclusivity Wang and Yao (2024). In Fan et al. (2024a), a relation between QET and quantum steering was suggested.
@@
As research on QET has progressed, a deeper understanding of the protocol has been achieved. It has been argued that the energy extracted by the receiver is not physically transported from the sender. Instead, the sender’s classical message enables the receiver to access energy that was previously unavailable. This argument was supported in Wu et al. (2024), where the receiver’s extracted energy can exceed the sender’s injected energy, indicating that the extracted energy cannot originate from the sender. In this sense, the QET protocol does not truly teleport energy but rather activates energy that was inaccessible to the receiver prior to the implementation of the protocol Rodríguez-Briones et al. (2023). This inaccessibility of energy extraction is rigorously described by the concept of strong local passive (SLP) states Frey et al. (2014); Alhambra et al. (2019). An SLP state is a multipartite quantum state, denoted by $\rho$, where no local quantum operation $\mathcal{G}_{B}$ applied to a subsystem can extract energy from the total system, characterized by the Hamiltonian $H$. Mathematically, this is expressed as the nonnegativity of the energy difference:
@@
In a QET protocol, the two parties share an SLP state, preventing the receiver from locally extracting energy. The sender then measures its local qubit, but the two-qubit system remains in an SLP state, meaning the receiver still cannot extract energy. However, classical communication between the sender and receiver breaks this limitation, making $\Delta E$ negative. A negative $\Delta E$ indicates that the receiver can locally extract energy, overcoming the restriction imposed by strong local passivity.
However, as has been noted, the QET protocol has traditionally been thought to impose strict constraints on the initial shared multipartite state and the observables employed by the sender during measurement. In this work, we challenge this notion and, for the first time, demonstrate that these constraints are not necessary. Specifically, we propose a novel QET protocol that relaxes these restrictions by introducing the concept of a local “effective Hamiltonian.” The relaxation of these constraints reveals that scenarios in which the energy within a system remains locally inaccessible are more frequent than previously believed. Additionally, the introduction of the local effective Hamiltonian simplifies the identification of optimal operations for the receiver to extract maximal energy, eliminating the need for cumbersome optimization techniques. Therefore, our results not only deepen the understanding of QET and its connection to SLP states but also expand its potential applications in quantum science and technology.
Finally, we validate our new protocol on quantum hardware, confirming both their theoretical correctness and experimental feasibility.
II Quantum energy teleportation and locally inaccessible eneergy
To begin, we revisit the minimal model of QET introduced in Hotta (2010c), which involves two qubits, $A$ and $B$, shared between the sender (Alice) and the receiver (Bob). The Hamiltonian of the system is given by:
@@ II Quantum energy teleportation and locally inaccessible eneergy
where $h$ and $\kappa$ are positive constants, and $X_{i}$ and $Z_{i}$ (with $i\in\{A,B\}$) are the Pauli-$X$ and Pauli-$Z$ operators for qubit $i$. The two qubits are initially prepared in the ground state $|g\rangle$ of $H_{AB}$, expressed as:
@@
$[\ |+\rangle\langle+|,2\kappa X_{A}\otimes X_{B}\ ]=[\ |-\rangle\langle-|,2\kappa X_{A}\otimes X_{B}\ ]=0.$
This commutativity ensures that Alice’s measurement, while injecting energy into the two-qubit system and altering its state away from the ground state $|g\rangle$, does not disturb the terms associated with Bob’s subsystem. Specifically, the measurement only affects the expectation value of the $-hZ_{A}$ term in Eq. (2), which corresponds to Alice’s qubit. In contrast, the terms associated with Bob, $-hZ_{B}+2\kappa X_{A}\otimes X_{B}$, remain unaffected, with their expectation values preserved as in the ground state $|g\rangle$. This guarantees that any energy extracted by Bob originates entirely from the previously inaccessible energy within the ground state $|g\rangle$, rather than from the energy injected by Alice.
Notably, after Alice’s measurement, the state of the system becomes:
@@
For Bob to access the inaccessible energy within Eq. (4), the QET protocol requires Alice to communicate her measurement outcome to Bob, based on which, Bob applies a conditional operator
@@
to his qubit, where $\tan(2\phi)\equiv 2\kappa/h$, with the sign $\pm$ depending on Alice’s measurement outcome. $Y$ is the Pauli-Y operator. The resulting state, expressed as:
@@
$\Delta E=\text{Tr}(H_{AB}\rho_{f})-\text{Tr}(H_{AB}\rho_{\text{SLP}})=-2h\sin^{2}(\phi-\theta)<0.$ (7)
This negativity indicates that Bob’s conditional operation successfully breaks the limit of strong local passivity, allowing him to extract the previously inaccessible energy that would remain unavailable without Alice’s classical communication.
Since the initial discovery of QET, the joint system has always been prepared in the ground state $|g\rangle$ of a given Hamiltonian, corresponding to the lowest possible energy configuration where both Alice’s and Bob’s qubits have minimal energy. When Alice measures her qubit, energy is injected into the system, deviating the state from $|g\rangle$. However, this energy remains localized to Alice’s subsystem because the measurement observable commutes with the interaction Hamiltonian. As a result, Bob’s local energy remains unchanged at its minimum value, preventing him from extracting energy immediately after Alice’s measurement.
Once Alice communicates her measurement outcome to Bob, he performs the conditional operation $G_{\pm}$ to extract energy. Since Bob’s local energy is kept at its minimum, the energy he extracts was believed to be transported from the energy injected by Alice during her measurement. This understanding gives rise to the term “quantum energy teleportation” (QET). Importantly, QET requires classical communication between Alice and Bob, ensuring that energy transfer does not exceed the speed of light, thus preserving causality Diener (1997). This is conceptually analogous to QST, where classical communication similarly ensures causality, preventing superluminal transfer of quantum information.
However, it was later realized that this interpretation of QET was incorrect Rodríguez-Briones et al. (2023). In fact, energy is not transported during the QET process. Although $|g\rangle$ is the ground state of the total Hamiltonian $H_{AB}$, it is not the ground state of the partial Hamiltonian, $-hZ_{B}+2\kappa X_{A}\otimes X_{B}$, obtained by excluding terms that do not involve Bob. This means that there is potential for Bob to extract energy from the system by locally driving his qubit to the ground state of this partial Hamiltonian. However, this extraction is prevented because the joint state remains an SLP state both before and after Alice’s measurement. In this framework, the classical communication in the QET protocol does not physically transfer energy but instead enables Bob to unlock and access the previously inaccessible energy stored within his subsystem Wu et al. (2024). The conclusion that no energy is transferred during QET can be further confirmed by the fact that Alice’s local energy, the expectation value of $-hZ_{A}$, remains unaffected by Bob’s conditional operations.
Although QET has been discussed in a close association with SLP states, the protocol remains widely misunderstood. To date, three constraints have been considered necessary for any QET protocol, but we demonstrate that none of these conditions are strictly required:
@@
    Ground State: It has been argued that if a density matrix $\rho$ commutes with the Hamiltonian, $[\rho,H_{AB}]=0$, and the population of the ground state $|g\rangle$ exceeds a threshold $p^{*}$ (as defined in Frey et al. (2014)), then $\rho$ is an SLP state. Consequently, all existing QET protocols initialize the system in the ground state $|g\rangle$ of $H_{AB}$, ensuring the state after Alice’s measurement retains sufficient population in $|g\rangle$ to preserve the SLP property. However, this condition is only sufficient, not necessary. We show that QET protocols can function even when the system begins in an excited pure state with zero population in $|g\rangle$. After Alice’s measurement, the resulting state can still be an SLP state, requiring Bob to rely on Alice’s information for energy extraction.
@@
    Commutativity: It is widely assumed that the observable used by Alice for her measurement must commute with the interaction term in $H_{AB}$ to prevent her measurement from injecting energy into Bob’s surroundings, hence preserving the SLP property. However, this condition is not strictly necessary. We show that even when Alice’s measurement observable does not commute with the interaction term, the interaction Hamiltonian can remain unaffected. Additionally, the resulting state still retains its SLP property.
@@
    Entanglement: Finally, it is believed that the shared state between Alice and Bob must be entangled for Alice’s measurement to have any influence on Bob’s subsystem. Relations were studied between the entanglement breaking due to Alice’s measurement and teleported energy Hotta (2010c). Arguments also suggested that if the ground state is entangled and Alice’s measurement disentangles the system, it becomes impossible for Bob, using only local operations and classical communications, to restore the entangled ground state and extract energy Alhambra et al. (2019). Along the same line, it was also suggested that quantum resources can improve the energy extraction efficiency for QET Fan et al. (2024b). However, we demonstrate that entanglement is not a strict requirement for QET. By introducing the concept of a local effective Hamiltonian, we show that Alice’s actions can influence Bob’s subsystem not only by changing the “departure state” through entanglement but also by changing the “destination state” via the effective Hamiltonian. Thus, the requirement of entanglement can be relaxed.
III A QET protocol relaxing all three constraints
@@
Compared to Eq. (2), this Hamiltonian includes an additional $Y_{A}\otimes Y_{B}$ interaction term. Consequently, the total interaction term, $(X_{A}\otimes X_{B}+Y_{A}\otimes Y_{B})$, cannot be factorized.
We consider the deep strong coupling regime Forn-Díaz et al. (2019), setting $\kappa>h>0$. In this regime, the Hamiltonian in Eq. (8) has four eigenstates with corresponding eigenvalues:
@@
To verify whether this state is SLP for Bob according to the Hamiltonian Eq. (8), one would need to check the inequality in Eq. (1), which involves parameterizing Bob’s general operation with 12 parameters for a qubit completely positive trace-preserving (CPTP) map Nielsen and Chuang (2010); Boyd and Vandenberghe (2004); Watrous (2018); Sim et al. (2020). However, a necessary and sufficient condition for determining whether a state $\rho$ is SLP for a Hamiltonian $H$ is given by a $4\times 4$ matrix $M(\rho,H)$, the form of which is defined in Alhambra et al. (2019) and given in Appendix A. Specifically, $\rho$ is SLP for $H$ if and only if $M(\rho,H)$ is positive semi-definite.
@@
In our case, it is straightforward to verify that both $M(|00\rangle\langle 00|,H_{AB})$ and $M(\rho_{\text{SLP}},H_{AB})$ share the identical eigenvalue set, $\{2,0,0,0\}$, all of which are nonnegative, confirming that these two matrices are positive semi-definite. Thus, starting from the first excited state $|00\rangle$, Bob cannot extract energy locally. After Alice’s measurement, the state evolves to Eq. (10), and Bob still cannot extract energy locally. Instead, Bob needs communication from Alice for his energy extraction.
@@
In contrast, if we had started with any of the other three eigenstates in Eq. (9), including the ground state $(|01\rangle-|10\rangle)/\sqrt{2}$, the resulting states after Alice’s measurement would not maintain the SLP property. Specifically, starting from the ground state, and performing Alice’s measurement, we compute the $M$ matrix of the resulting state and find its four eigenvalues to be $\{2,1.5,-0.5,0\}$. This confirms that the post-measurement state is not an SLP state, meaning QET protocols are not necessary for Bob’s energy extraction. Therefore, our initial choice of the first excited state $|00\rangle$ is well justified.
@@
Through the new protocol, we observe that all three previously considered constraints for QET are relaxed. First, the system is initialized in an excited state where the ground-state population is zero. Second, although Alice’s measurement does not commute with the interaction term $Y_{A}\otimes Y_{B}$, it does not inject energy to this term, as $\langle Y_{A}\otimes Y_{B}\rangle$ remains 0 throughout the entire process. Finally, the system begins in the product state $|00\rangle$, meaning Alice’s measurement cannot directly affect Bob’s local state. However, for Bob to extract energy based on Alice’s measurement outcome, her action must influence Bob’s subsystem. To address this, we introduce the concept of a local effective Hamiltonian for Bob and show that Alice’s measurement modifies Bob’s local effective Hamiltonian rather than his local state.
@@
Since the two qubits remain disentangled after Alice’s measurement, the total energy of the system is given by
@@
where $\langle\cdot\rangle_{A}\equiv\langle\psi_{A}|\cdot|\psi_{A}\rangle$ denotes the expectation value of the operator $\cdot$ with respect to Alice’s state $|\psi_{A}\rangle$, and similar for $\langle\cdot\rangle_{B}$ with respect to Bob’s state.
If Alice’s measurement yields the outcome $+$, her state collapses to |+ket|$\rangle$| + ⟩, and the total energy becomes
@@
implying that Bob’s effective Hamiltonian is $-hZ_{B}+\kappa X_{B}$. Similarly, if Alice’s outcome is --, Bob’s effective Hamiltonian becomes $-hZ_{B}-\kappa X_{B}$.
@@
By assuming that Alice’s measurement outcome is $+$ (the analysis for -- follows similarly), the general two-qubit state is given by:
@@
Under Schrödinger’s equation, the time evolution of the state is:
@@
Using $\langle+|X|+\rangle=1$ and $\langle+|Y|+\rangle=\langle+|Z|+\rangle=0$, the evolution simplifies to:
@@
Similarly, for the -- outcome, Bob’s effective Hamiltonian becomes $-hZ_{B}-\kappa X_{B}$. Thus, the local effective Hamiltonian for Bob is:
@@
Figure 1: The quantum circuit implemented on IBM quantum hardware for verifying the feasibility of our proposed QET protocol. Since the system begins in the first excited state of the Hamiltonian in Eq. (8), $|0\rangle\otimes|0\rangle$, no additional state preparation is required. Alice initiates the protocol by performing a measurement, which injects energy into the system. After receiving Alice’s measurement outcome, Bob applies a conditional operation to extract energy from the system. A final measurement is then performed to verify the energy distribution within the system.
For Bob to extract energy using Alice’s information, we diagonalize the effective Hamiltonian Eq. (17). The ground states $|g^{\pm}\rangle$ and excited states $|e^{\pm}\rangle$ of the effective Hamiltonian, together with their eigenvalues, are given by:

### v2

# Strong Local Passivity in Unconventional Scenarios: A New Protocol for Amplified Quantum Energy Teleportation

@@ Abstract
Quantum energy teleportation (QET) has been proposed to overcome the restrictions of strong local passivity (SLP) and to facilitate energy transfer in quantum systems. Traditionally, QET has only been considered under strict constraints, including the requirements that the initial state be the ground state of an interacting Hamiltonian, that Alice’s measurement commute with the interaction terms, and that entanglement be present. These constraints have significantly limited the broader applicability of QET protocols. In this work, we demonstrate that SLP can arise beyond these conventional constraints, establishing the necessity of QET in a wider range of scenarios for local energy extraction. This leads to a more flexible and generalized framework for QET. Furthermore, we introduce the concept of a “local effective Hamiltonian,” which eliminates the need for optimization techniques in determining Bob’s optimal energy extraction in QET protocols. As an additional advantage, the amount of energy that can be extracted using our new protocol is amplified to be 7.2 times higher than that of the original protocol. These advancements enhance our understanding of QET and extend its broader applications to quantum technologies. To support our findings, we implement the protocol on quantum hardware, confirming its theoretical validity and experimental feasibility.
@@ I Introduction
In contrast, a related protocol, quantum energy teleportation (QET), was believed to enable the teleportation of energy Hotta (2009, ). Similarly to QST, QET relies on classical communication. However, QET protocols impose stricter constraints compared to QST. While QST only requires the sender and receiver to share an entangled state, QET further requires that this state be the ground state of an interacting Hamiltonian. Additionally, in QST, the sender performs a measurement and transmits the outcome to the receiver via a classical communication channel. In QET, however, the observable used for the measurement must also commute with the interaction term in the Hamiltonian, ensuring that the local energy of the receiver remains undisturbed prior to the teleportation process.
@@
Since its initial discovery in spin-chain systems Hotta (2008, 2009), QET has been extensively explored in many other physical systems. Theoretical studies have investigated its application in relativistic quantum field theory Hotta (2008); Ikeda (2023), trapped ions Hotta (2009), harmonic oscillators Hotta (2010), black-hole physics Hotta (2010), linear harmonic chains Nambu and Hotta (2010), quantum Hall systems Yusa et al. (2011), Gibbs spin particles Frey et al. (2013); Trevison and Hotta (2015), squeezed vacuum states Hotta et al. (2014), topological orders Ikeda (2023), and quantum information science Ikeda and Lowe (2024). More recently, experimental efforts have demonstrated progress toward realizing QET, including in an NMR system Rodríguez-Briones et al. (2023) and on superconducting quantum hardware Ikeda (2023). The combination of QST and QET was proposed to enable a long-range QET protocol Ikeda (2023). A trade-off relationship between QST and QET was proposed, suggesting their potential exclusivity Wang and Yao (2024). In Fan et al. (2024), a relation between QET and quantum steering was suggested.
@@
However, despite its name, the QET protocol does not physically teleport energy between distant locations. Instead, the sender’s classical message only enables the receiver to access energy that was previously inaccessible prior to the implementation of the protocol Rodríguez-Briones et al. (2023). This argument can be further supported in Wu et al. (2024), where the receiver’s extracted energy can exceed the sender’s injected energy, indicating that the extracted energy cannot originate from the sender. This inaccessibility of energy extraction can be rigorously described by the concept of strong local passive (SLP) states Frey et al. (2014); Alhambra et al. (2019). An SLP state is a multipartite quantum state, denoted by $\rho$, where no local quantum operation $\mathcal{G}_{B}$ applied to a subsystem can extract energy from the total system, characterized by the Hamiltonian $H$. Mathematically, this is expressed as the nonnegativity of the energy difference:
@@
In a QET protocol, the two parties share an SLP state, preventing the receiver from locally extracting energy. The sender then measures its local qubit, but the two-qubit system remains in an SLP state, meaning the receiver still cannot extract energy. However, classical communication between the sender and receiver breaks this limitation, making $\Delta E$ negative. A negative $\Delta E$ indicates that the receiver can locally extract energy, overcoming the restriction imposed by SLP.
However, previous QET protocols have been studied only under scenarios with strict constraints on the initial shared state and/or on the sender’s measurement observables. In our view, these constraints arise from the initial misconception—reinforced by the name “quantum energy teleportation”—that QET directly teleports energy between distant locations, as well as from insufficient attention to the connection between QET and SLP.
In this work, we demonstrate that SLP can arise beyond these conventional constraints, establishing the necessity of QET in a wider range of scenarios for local energy extraction. Specifically, by introducing the concept of a “local effective Hamiltonian,” we construct an SLP state that satisfies none of these restrictions, yet still forbids local energy extraction. As a result, energy extraction is only possible through a QET protocol. In this sense, relaxing these constraints reveals that situations where energy remains locally inaccessible are more common than previously thought, broadening the applicability of QET. As a separate and technical advantage, the local effective Hamiltonian simplifies the identification of optimal operations for the receiver, enabling maximal energy extraction without the need for cumbersome optimization techniques. Therefore, our results not only deepen the understanding of QET and its connection to SLP states but also expand its potential applications in quantum science and technology.
Finally, we implement our new protocol on quantum hardware, confirming both its theoretical correctness and experimental feasibility.
II Quantum Energy Teleportation and Locally Inaccessible Energy
To begin, we revisit the minimal model of QET introduced in Hotta (2010), which involves two qubits, $A$ and $B$, shared between the sender (Alice) and the receiver (Bob). The Hamiltonian of the system is given by
@@ II Quantum Energy Teleportation and Locally Inaccessible Energy
where $h$ and $\kappa$ are positive constants, and $X_{i}$ and $Z_{i}$ (with $i\in\{A,B\}$) are the Pauli-$X$ and Pauli-$Z$ operators for qubit $i$. The two qubits are initially prepared in the ground state $|g\rangle$ of $H_{AB}$, expressed as
@@
$|+\rangle\langle+|,2\kappa X_{A}\otimes X_{B}\ ]=[\ |-\rangle\langle-|,2\kappa X_{A}\otimes X_{B}\ ]=0.$
This commutativity ensures that Alice’s injected energy stays localized with her. Specifically, the measurement only affects the expectation value of the $hZ_{A}$ term in Equation (2), which corresponds to Alice’s qubit. In contrast, the terms associated with Bob, $hZ_{B}+2\kappa X_{A}\otimes X_{B}$, remain unaffected.
Notably, after Alice’s measurement, the state of the system becomes
@@
For Bob to access the inaccessible energy within Equation (4), the QET protocol requires Alice to communicate her measurement outcome to Bob, based on which Bob applies a conditional operator,
@@
to his qubit, where $\tan(2\phi)\equiv 2\kappa/h$, with the sign $\pm$ depending on Alice’s measurement outcome. $Y$ is the Pauli-Y operator. The resulting state, expressed as
@@
$\Delta E=\text{Tr}(H_{AB}\rho_{f})-\text{Tr}(H_{AB}\rho_{\text{SLP}})=-2\sin^{2}(\phi-\theta)\sqrt{h^{2}+4\kappa^{2}}<0.$ (7)
This negativity indicates that Bob’s conditional operation successfully breaks the limit of strong local passivity, allowing him to extract the previously inaccessible energy that would remain unavailable without Alice’s classical communication. For our parameter setting, $h=1$ and $\kappa=1.5$, which will be used in the next section, the extracted energy for this protocol is 0.1114.
It is important to clarify that after Alice’s measurement, the SLP property of the system is fragile and can vanish as the system evolves under $H_{AB}$. Consequently, Bob’s local energy extraction is, in principle, possible even without Alice’s message, provided he waits for an appropriate duration for the system to evolve and then applies a final local operation to extract energy. Nevertheless, we emphasize that the key advantage of QET lies in its speed. When relativistic constraints are not considered, Alice’s message can travel much faster than the natural timescale of the system’s evolution Diener (1997). This enables Bob to extract energy almost instantaneously, distinguishing QET from conventional energy extraction mechanisms.
We point out that the following constraints have been widely assumed for previous QET protocols. However, these constraints are intrinsically irrelevant to the SLP property of quantum states and should not be assumed as fundamental for applications of QET. Yet, they still appear consistently in discussions of QET. In this work, we demonstrate that these constraints are not required for a state to exhibit SLP. Consequently, even when these constraints are relaxed, QET remains essential for enabling local energy extraction. The constraints are as follows:
@@
    Ground State: It has been argued that if a density matrix $\rho$ commutes with the Hamiltonian, $\rho,H_{AB}]=0$, and the population of the ground state $|g\rangle$ exceeds a threshold $p^{*}$ (as defined in Frey et al. (2014)), then $\rho$ is an SLP state. Consequently, most existing QET protocols initialize the system in the ground state $|g\rangle$ of $H_{AB}$, ensuring that the state after Alice’s measurement retains sufficient population in $|g\rangle$ to preserve the SLP property. However, this condition is only sufficient, not necessary. We show that a system can exhibit SLP even when initialized in an excited pure state with zero population in $|g\rangle$. Moreover, after Alice’s measurement, the resulting state can still be an SLP state, requiring Bob to rely on Alice’s information for energy extraction. While it is intuitive that energy extraction is forbidden from a ground state, it is counter-intuitive that energy can also be blocked for certain excited states. This suggests that the common understanding that only ground states prevent energy extraction may be incomplete.
@@
    Commutativity: To construct an SLP scenario that prevents Bob from locally extracting energy, it is typically assumed that Alice’s measurement does not increase the energy in Bob’s surroundings, which further requires Alice’s measurement observable to commute with the interaction term in $H_{AB}$, ensuring that her measurement does not inject energy into Bob’s surroundings. However, we emphasize that the SLP property stands on its own and is independent of this assumption. We demonstrate that even when Alice’s measurement observable does not commute with the interaction term, the resulting state can still exhibit SLP. Consequently, Bob still requires the QET protocol for local energy extraction.
@@
    Entanglement: It was believed that the shared state between Alice and Bob must be entangled for Alice’s measurement to have any influence on Bob’s subsystem. It was believed that “a large amount of teleported energy requests a large amount of consumption of the ground-state entanglement between A and B in this model” Hotta . Relations were further studied between the entanglement breaking due to Alice’s measurement and the teleported energy Hotta (2010). Arguments also suggested that if the ground state is entangled and Alice’s measurement disentangles the system, it becomes impossible for Bob, using only local operations and classical communications, to restore the entangled ground state and extract energy Alhambra et al. (2019). Along the same line, it was also suggested that quantum resources can improve the energy extraction efficiency for QET Fan et al. (2024). However, we demonstrate that entanglement is not a strict requirement for the exhibition of SLP and the application of QET. In fact, in our proposed protocol, the shared state between Alice and Bob remains disentangled throughout the entire process, yet the system still exhibits the SLP property. To understand this, we introduce the concept of a local effective Hamiltonian. We show that Alice’s actions can influence Bob’s subsystem not only by changing the “departure state” through entanglement but also by changing the “destination state” via Bob’s local effective Hamiltonian. As long as either the departure or destination state is affected, conditional operations are required to break SLP for Bob’s local energy extraction. Therefore, the necessity of entanglement can be relaxed if Alice’s measurements can influence the destination state alone.
We note that some previous works have gone beyond one or two of the constraints discussed above. For example, Ref. Frey et al. (2013) considered a QET protocol with an initial Gibbs state that may contain no entanglement, while Ref. Wu et al. (2024) introduced a class of initial states that are not ground states. However, in both cases, the SLP condition was analyzed only before Alice’s measurement—not afterward. It can be readily verified that their post-measurement states can violate the SLP condition, meaning that classical communication is not required for Bob’s energy extraction. As a result, these protocols do not truly require the implementation of QET. In contrast, our work not only emphasizes the importance of verifying the SLP condition after Alice’s measurement but also relaxes all three constraints within one single and minimal protocol.
III Strong Local Passivity Relaxing Previous Constraints
@@
Compared to Equation (2), this Hamiltonian replaces one of the two $X_{A}\otimes X_{B}$ terms with a $Y_{A}\otimes Y_{B}$ interaction term. Consequently, the total interaction term, $(X_{A}\otimes X_{B}+Y_{A}\otimes Y_{B})$, cannot be factorized. As we will show later, this replacement will greatly enhance the amount of energy that can be extracted by Bob.
We consider the deep strong coupling regime Forn-Díaz et al. (2019), setting $\kappa>h>0$. In this regime, the Hamiltonian in Equation (8) has four eigenstates with corresponding eigenvalues:
@@
To verify whether this state is SLP for Bob according to the Hamiltonian Equation (8), one would need to check the inequality in Equation (1), which involves parameterizing Bob’s general operation with 12 parameters for a qubit completely positive trace-preserving (CPTP) map Nielsen and Chuang (2010); Boyd and Vandenberghe (2004); Watrous (2018); Sim et al. (2020). However, a necessary and sufficient condition for determining whether a state $\rho$ is SLP for a Hamiltonian $H$ is given by a $4\times 4$ matrix $M(\rho,H)$, the form of which is defined in Alhambra et al. (2019) and given in Appendix A. Specifically, $\rho$ is SLP for $H$ if and only if $M(\rho,H)$ is positive semi-definite.
@@
In our case, it is straightforward to verify that both $M(|00\rangle\langle 00|,H_{AB})$ and $M(\rho_{\text{SLP}},H_{AB})$ share the identical eigenvalue set, $\{2,0,0,0\}$, all of which are nonnegative, confirming that these two matrices are positive semi-definite. Thus, starting from the first excited state $|00\rangle$, Bob cannot extract energy locally. After Alice’s measurement, the state evolves to Equation (10), and Bob still cannot extract energy locally. Instead, Bob needs communication from Alice for his energy extraction.
@@
In contrast, if we had started with any of the other three eigenstates in Equation (9), including the ground state $(|01\rangle-|10\rangle)/\sqrt{2}$, the resulting states after Alice’s measurement would not maintain the SLP property. Specifically, starting from the ground state, and performing Alice’s measurement, we compute the $M$ matrix of the resulting state and find its four eigenvalues to be $\{2,1.5,-0.5,0\}$. This confirms that the post-measurement state is not an SLP state, meaning QET protocols are not necessary for Bob’s energy extraction. Therefore, our initial choice of the first excited state $|00\rangle$ is well justified.
@@
Through the new protocol, we observe that all three previously considered constraints for QET are relaxed. First, the system is initialized in an excited state where the ground-state population is zero. Second, although Alice’s measurement does not commute with the interaction term $Y_{A}\otimes Y_{B}$, it does not inject energy into this term, as $\langle Y_{A}\otimes Y_{B}\rangle$ remains 0 throughout the entire process. Finally, the system begins in the product state $|00\rangle$, meaning Alice’s measurement cannot directly affect Bob’s local state.
At first glance, it may seem puzzling that Alice’s measurement is needed in a product-state construction, as the measurement cannot affect Bob’s local state. To address this, below we introduce the concept of a local effective Hamiltonian for Bob. We show that, for energy extraction, Bob must rotate his local state to the ground state of this effective Hamiltonian. In a product-state setup, Alice’s measurement does not alter Bob’s initial (“departure”) state, but it does modify the effective Hamiltonian—and thus the ground (“destination”) state of this Hamiltonian that Bob must reach. As a result, classical communication from Alice remains necessary to inform Bob of the appropriate rotation. In general, conditional operations are required to break the SLP constraint as long as one of the departure or destination states is contingent on Alice’s measurement. Notably, this allows the relaxation of the entanglement requirement when Alice’s measurement only influences the destination state.
@@
Now we introduce Bob’s local effective Hamiltonian. Since the two qubits remain disentangled after Alice’s measurement, the total energy of the system is given by
@@
where $\langle\cdot\rangle_{A}\equiv\langle\psi_{A}|\cdot|\psi_{A}\rangle$ denotes the expectation value of the operator $\cdot$ with respect to Alice’s state $|\psi_{A}\rangle$, and similarly for $\langle\cdot\rangle_{B}$ with respect to Bob’s state.
If Alice’s measurement yields the outcome ++, her state collapses to $|+\rangle$, and the total energy becomes
@@
implying that Bob’s effective Hamiltonian is $hZ_{B}+\kappa X_{B}$. Similarly, if Alice’s outcome is -, Bob’s effective Hamiltonian becomes $hZ_{B}-\kappa X_{B}$.
@@
By assuming that Alice’s measurement outcome is ++ (the analysis for - follows similarly), the general two-qubit state is given by
@@
Under Schrödinger’s equation, the time evolution of the state is
@@
Using $\langle+|X|+\rangle=1$ and $\langle+|Y|+\rangle=\langle+|Z|+\rangle=0$, the evolution simplifies to
@@
Similarly, for the - outcome, Bob’s effective Hamiltonian becomes $hZ_{B}-\kappa X_{B}$. Thus, the local effective Hamiltonian for Bob is
@@
For Bob to extract energy using Alice’s information, we diagonalize the effective Hamiltonian Equation (17). The ground states $|g^{\pm}\rangle$ and excited states $|e^{\pm}\rangle$ of the effective Hamiltonian, together with their eigenvalues, are given by